OpenStudy (anonymous):

Suppose you invest $2500 at an annual interest rate of 3% compounded continuously. How much will you have in the account after 7 years? Round the solution to the nearest dollar. The steps for this would be great, thanks!

OpenStudy (anonymous):

\[A = Pe^{rt}\]

OpenStudy (anonymous):

Are you familiar with this formula?

OpenStudy (anonymous):

No :/

OpenStudy (anonymous):

actually it sounds familiar but I don't know it.

OpenStudy (anonymous):

Alright. This is the formula for compounded continuously. Where: P = 2500 e = 2.718... r = 0.03 t = 7 Just plug in numbers :)

OpenStudy (anonymous):

awesome. Thank you!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

So then, the answer is....20411?

OpenStudy (anonymous):

nope I did something wrong

OpenStudy (anonymous):

I got 111,582,957

OpenStudy (anonymous):

oh! i used .3 instead of .03. my mistake.

OpenStudy (anonymous):

\[A = 2500(2.718...)^{0.03(7)}\]\[A = 2500(2.718)^{0.21}\]\[A = 2500(1.23367806)\]\[A = 3084.20\]

OpenStudy (anonymous):

Ohhh, you multiply the 2500 last. Alright, thanks a lot.

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

Wait how did you get all of those numbers for Pe^rt?